OpenStudy (anonymous):

If f is a function such that the limit as x approaches a of the quotient of the quantity f of x minus f of a and the quantity x minus a equals 5, then which of the following statements must be true? f(a) = 5 The slope of the tangent line to the function at x = a is 5. The slope of the secant line through the function at x = a is 5. The linear approximation for f(x) at x = a is y = 5

OpenStudy (anonymous):

Will medal and fan!

OpenStudy (freckles):

\[\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}=5\] is the given right

OpenStudy (freckles):

I will give you a hint isn't the left hand side the definition of a magical d word.

OpenStudy (anonymous):

The answer is the second one, I believe.

OpenStudy (freckles):

yeah \[\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}=f'(a)\] and we are given f'(a)=5

OpenStudy (freckles):

f'(a)=5 means f has slope at x=a as 5

OpenStudy (freckles):

very good

OpenStudy (anonymous):

Thank you!

OpenStudy (jhannybean):

It's not the slope of the secant line, btw.

OpenStudy (freckles):

The second one is " The slope of the tangent line to the function at x = a is 5. "

OpenStudy (freckles):

That is how I read the choices

OpenStudy (freckles):

f(a)=5 is the first choice right?

OpenStudy (anonymous):

What @freckles said is correct. That is the correct answer also.

OpenStudy (jhannybean):

Oh is that really the second one??? I thought f(a) = 5 was part of the question xD Lmao, sorry then

OpenStudy (jhannybean):

Oh that's what I was confused about earlier lol.

OpenStudy (freckles):

yeah we don't have two points so it can't be secant line

OpenStudy (freckles):

I'm just saying this for clarity purposes

OpenStudy (freckles):

\[\frac{f(x)-f(a)}{x-a} \text{ would be slope of the secant line through points } \\ (x,f(x)) \text{and } (a,f(a)) \\ \lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a} \text{ would be slope of tangent line at } x=a \]

OpenStudy (jhannybean):

Perfect!